OpenStudy (anonymous):

A=1/7h(x+y). x=? Please show with steps how you arrive at the answer.

OpenStudy (anonymous):

Well is this it? \[A=\frac{1}{7}h(x+y)\]

OpenStudy (anonymous):

Yes.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

So then try this. To get rid of the \(\frac{1}{7}\), Multiply by 7: \[7A=h(x+y)\] Then divide both sides by \(h\): \[\frac{7A}{h}=x+y\] Then subtract both sides by y: \[\frac{7A}{h}-y=x\]

OpenStudy (anonymous):

Maybe I'm viewing this wrong. But I know the answer is 7A-hy/h. That's not what you have for an answer.